# System Prompt Optimization with Meta-Learning

### v1

@@ 1 Introduction
However, despite various studies on prompt optimization, they have mainly focused on user prompts, while overlooking the system prompts that are equally important components in shaping the behavior of LLMs and additionally have the potential to steer their responses across far more diverse contexts. Notably, there are a couple of benefits in optimizing system prompts. First, system prompts serve as the foundational instructions that are invariant, meaning that a single well-optimized system prompt can generalize across multiple tasks and domains. Second, the optimized system prompts can establish a robust behavioral framework, enabling LLMs to adapt more robustly to unseen user prompts and domains, while having the potential to create a synergistic relationship with user prompts.
@@ 2 Related Work
Figure 2: Overview of MetaSPO, which consists of the inner loop for user prompt optimization and the outer loop for system prompt optimization, operationalized through the meta-learning framework. (A) Inner Loop generates candidate user prompts by analyzing incorrectly predicted examples, and then evaluates them with the system prompt to select refined prompts for individual tasks. (B) Outer Loop generates candidate system prompts by analyzing incorrect examples from all source tasks, and then evaluates them across various user prompts and tasks to ensure generalizability.
@@ 3.1 Preliminary
Given the sensitivity of LLMs to the prompts they receive, prompt optimization emerges as an effective solution, whose objective is to automatically discover prompts that maximize task performance through iterative refinement processes, guided by performance feedback from available examples. Formally, let ${T}$ denote a task, which represents a dataset or distribution of input-output pairs $({\bm{q}},{\bm{a}})$, where ${\bm{q}}$ is the query and ${\bm{a}}$ is the ground truth answer. Then, the goal of prompt optimization is to identify the optimal prompt ${\bm{u}}^{*}$ (starting from the initial prompt ${\bm{u}}$) that maximizes the task performance on ${T}$. To achieve this, previous works [45, 27, 37] typically formulate the objective function (and propose various methods to optimize it), as follows:
@@ 3.3 MetaSPO: Meta-level System Prompt Optimizer
It is worth noting that since MetaSPO is designed as a general framework, it allows the use of any off-the-shelf prompt optimization techniques, and, among many, one instantiation is to iteratively update the prompt to correct examples that were previously handled incorrectly, thereby improving the overall performance on the target tasks, which is similar to Pryzant et al. [27]. Specifically, as the first step, we measure the performance of the current user prompt on examples from the target task, and identify responses that are incorrect. To improve the performance, we then aim to refine the prompt to address these errors, and, to achieve this, we conduct a failure analysis, where we prompt the LLM with the current user prompt and incorrect examples to uncover the underlying issues. After that, based on the user prompt and its analyzed problems, we further prompt the LLM multiple times to generate (potentially refined) candidate user prompts. However, as not all prompts generated result in performance improvement, we measure the performance of prompts (including previously used ones) on the target task, and select the top $k$ prompts that perform the best.
@@ 4.1 Experimental Setup
To evaluate MetaSPO, we consider two scenarios that reflect real-world applications of the optimized system prompt. First, in the Unseen Generalization scenario where there is no data available for the target task, the optimized system prompt is directly applied to test-time queries and user prompts without any additional optimization on them. Meanwhile, in the Test-Time Adaptation scenario, the user prompts provided are further optimized using a small number of examples from the target task, while the optimized system prompt remains fixed. In addition to them, we further consider two different settings in optimizing system prompts: Global where the goal is to obtain the global system prompt that is generalizable across domains, and Domain, a more relaxed case where the system prompt is designed and deployed to handle tasks and queries within one specific domain [^1] Unless otherwise stated, we report results with the Domain setup..
@@
Default 45.1 68.9 64.0 59.9 17.5 52.4
SPRIG 45.4 69.3 65.3 64.7 17.7 53.6
MetaSPO 45.6 71.4 67.3 67.2 19.9 55.2
Figure 6: Efficiency for test-time adaptation as the function of optimization iterations (left) and data quantity (right). The results are averaged over Review Analysis and Reasoning domains. The gray dashed line indicates the final performance of the Default baseline.
@@ 4.2 Experimental Results and Analyses
We first report the performance of optimized system prompts on the unseen generalization scenario, where examples for target tasks are not available for prompt optimization; thus, the user prompts are not optimized on them, and, for evaluation, (ten) user prompts on each task are obtained via LLM prompting (see A.5 for details). As shown in Table 1, we then observe that MetaSPO consistently outperforms all baselines across both the Global and Domain system prompt optimization settings. To see whether the optimized system prompt contributes to the performance gain when coupled with diverse user prompts, we visualize the performance of (randomly sampled 20%) user prompts using the system prompt optimized from MetaSPO, compared to the default system prompt in Figure 5 (see B.4 for the break-down results by domain). From this, we observe that 85.0% of the user prompts exhibit improved performance with MetaSPO, indicating that it effectively and robustly enhances the performance across a broad range of user prompts.
@@
We hypothesize that if the target tasks (used for evaluation) are more similar to the source tasks (used for meta-learning), the system prompt optimized from source tasks is more effective for target tasks. To confirm this, we measure the similarity of examples between source and target tasks using either Bag-of-Words rank correlation [21] or cosine similarity over the embedding space with the language model [33], then measure the Pearson correlation of the similarity with its corresponding performance improvement (over the Default system prompt). The results in Figure 5 show a positive correlation between the source-target task similarity and the improvement, with a Pearson correlation coefficient of +0.56. Also, the cosine similarity result in B.5 shows a positive correlation with performance improvement. Yet, more interestingly, we observe that MetaSPO remains effective even for low-similarity tasks, yielding performance gains. These results demonstrate that, while including more source tasks that are close to target tasks is beneficial, MetaSPO enables learning generalizable knowledge that is useful across diverse tasks.
@@
Table 3: Results with different LLMs for MetaSPO, where Cross Model refers to prompts optimized with Llama3.2 (3B) and applied to other models. Results are averaged over Review Analysis and Reasoning domains. Numbers in bold indicate the highest followed by underline.
@@
[…]
Table 4: Variations of MetaSPO, where we use different compositions of prompt optimizers within our meta-learning framework. We report the average performance score for each domain. The highest score in each domain is highlighted in bold, while the second-highest is underlined.
@@
We conduct an auxiliary analysis to examine whether MetaSPO is versatile with other LLMs (for response generation) and whether the optimized system prompt from one LLM can be generalizable to other LLMs. For both experiments, we consider the following LLMs: Llama 3.1 (8B), Qwen 2.5 (7B), and GPT-4o mini. We then report the results in subsection 4.2, and, from this, we observe that MetaSPO demonstrates its robustness and generalizability. Specifically, the system prompt optimized by MetaSPO is superior to other baselines regardless of the underlying LLMs used as the base model. Furthermore, in the Cross Model, the system prompt optimized for Llama 3.2 (3B) demonstrates strong generalization capabilities when applied to LLMs other than it without requiring additional optimization. Overall, these results confirm that MetaSPO is effective for diverse LLMs, producing system prompts that not only perform well within the LLM they were optimized for but also maintain high performance across different LLMs. Lastly, we further extend our analysis of MetaSPO with different optimizer LLMs and show its robustness, provided in B.6.
@@
We present examples for the optimized system prompts in Appendix C, from which we observe that they typically provide a more specific role to the LLM beyond the helpful assistant (e.g., you are a knowledgeable and analytical assistant designed to process and analyze information across a broad range of topics) and include detailed guidelines or objectives on top of it.
@@ 5 Conclusion
In this paper, we introduced the novel problem of bilevel system prompt optimization, which differs from existing prompt optimization work (that primarily targets optimization of task-specific user prompts) and instead focuses on designing system prompts that are robust to diverse user prompts and transferable to unseen tasks and domains. To address this problem, we then proposed MetaSPO, a meta-learning framework that optimizes system prompts over a wide range of user prompts and tasks, while iteratively refining the user prompts and using those optimized user prompts during system prompt optimization to ensure effective generalization of the system prompt to various (optimized) user prompts. We extensively validated the proposed MetaSPO framework on 14 unseen datasets across 5 distinct domains, on which we demonstrated that it consistently outperforms baselines in both unseen generalization and test-time adaptation scenarios. We believe our work establishes a significant step forward in enhancing the robustness and adaptability of LLMs, by enabling the use of the optimized system prompts designed to generalize across diverse tasks and domains as well as various LLMs. We discuss the limitations and societal impacts of our work in Appendix D.
@@ A.2 Additional Implementation Details
We now provide the additional implementation details for other optimization methods included in our experiments. Regarding ProTeGi [27] (presented in Table 6), it performs six iterations with a beam size of three. For the Outer Loop, we perform six iterations, which is twice the number of iterations used in MetaSPO. This adjustment ensures an equivalent total number of iterations, as the Outer Loop method lacks an inner loop process. In the case of MetaSPO w/ APE and MetaSPO w/EVO (that is operationalized without the prompt analysis step), we generate 18 new candidate prompts through resampling, crossover, and mutation. For SPRIG, we conduct the experiments using the implementation provided in its official repository referenced in the original paper [43]. We iterate SPRIG three times to ensure a comparable amount of computation to the proposed MetaSPO (See B.3 for details). Experiments are primarily conducted using an NVIDIA A5000 GPU.
@@ A.3 Algorithm of MetaSPO
0:  Task distribution $\mathcal{T}$, Initial system prompt ${\bm{s}}$, Number of iterations $N$
@@ A.5 Generation of User Prompts for Unseen Generalization
In the unseen generalization setting, we evaluate whether the system prompt is effective with user prompts not optimized for target tasks. To generate these unoptimized user prompts, we provide GPT-4o mini with ten input-output example pairs, generating the coarse user prompts following Zhou et al. [45] (please see Table 11 for the detailed meta prompt used to generate ten different user prompts). After that, the answer format prompts (e.g. At the end present your answer in <answer>yes</answer> or <answer>no</answer>.) are added to those generated user prompts. Note that the examples of the generated user prompts for each target task are provided in LABEL:tab:example_up_unseen_gen.
Table 11: Meta Prompt for generating ten different user prompts for the unseen generalization scenario.
@@
Table 12: User Prompts for Unseen Generalization Experiments, where we sample three among ten for each task.
Target Tasks User Prompts

### v2

@@ 1 Introduction
However, despite various studies on prompt optimization, they have mainly focused on user prompts for solving a specific task, while overlooking the system prompts that could significantly influence response generation and have far-reaching effects across diverse tasks and domains. Notably, there are a couple of benefits in optimizing system prompts. First, system prompts serve as the foundational instructions that are invariant, meaning that a single well-optimized system prompt can generalize across multiple tasks and domains. Second, the optimized system prompts can establish a robust behavioral framework, enabling LLMs to adapt more robustly to unseen user prompts and domains, while having the potential to create a synergistic relationship with user prompts.
@@ 2 Related Work
Figure 2: Overview of MetaSPO, which consists of the inner loop for user prompt optimization and the outer loop for system prompt optimization, operationalized through the meta-learning framework. (A) Inner Loop generates candidate user prompts by analyzing incorrectly predicted examples and then evaluates them with the system prompt to select refined prompts for individual tasks. (B) Outer Loop generates candidate system prompts by analyzing incorrect examples from all source tasks, and then evaluates them across various user prompts and tasks to ensure generalizability.
@@ 3.1 Preliminary
Given the sensitivity of LLMs to the prompts they receive, prompt optimization emerges as an effective solution, whose objective is to automatically discover prompts that maximize task performance through iterative refinement processes, guided by performance feedback from available examples. Formally, let T{T} denote a task, which represents a dataset or distribution of input-output pairs $({\bm{q}},{\bm{a}})$, where $\bm{q}$} is the query and $\bm{a}$} is the ground truth answer. Then, the goal of prompt optimization is to identify the optimal prompt $\bm{u}$$^{*}$ (starting from the initial prompt $\bm{u}$}) that maximizes the task performance on T{T}. To achieve this, previous works [45, 27, 37] typically formulate the objective function (and propose various methods to optimize it), as follows:
@@ 3.3 MetaSPO: Meta-level System Prompt Optimizer
It is worth noting that since MetaSPO is designed as a general framework, it allows the use of any off-the-shelf prompt optimization techniques, and, among many, one instantiation is to iteratively update the prompt to correct examples that were previously handled incorrectly, thereby improving the overall performance on the target tasks, which is similar to Pryzant et al. [27]. Specifically, as the first step, we measure the performance of the current user prompt on examples from the target task and identify responses that are incorrect. To improve the performance, we then aim to refine the prompt to address these errors, and to achieve this, we conduct a failure analysis, where we prompt the LLM with the current user prompt and incorrect examples to uncover the underlying issues. After that, based on the user prompt and its analyzed problems, we further prompt the LLM multiple times to generate (potentially refined) candidate user prompts. However, as not all generated prompts result in performance improvement, we measure the performance of prompts (including previously used ones) on the target task, and select the top $k$ prompts that perform the best.
@@ 4.1 Experimental Setup
To evaluate MetaSPO, we consider two scenarios that reflect real-world applications of the optimized system prompt. First, in the Unseen Generalization scenario, where there is no data available for the target task, the optimized system prompt is directly applied to test-time queries and user prompts without any additional optimization on them. Meanwhile, in the Test-Time Adaptation scenario, the user prompts provided are further optimized using a small number of examples from the target task, while the optimized system prompt remains fixed. In addition to them, we further consider two different settings in optimizing system prompts: Global where the goal is to obtain the global system prompt that is generalizable across domains, and Domain, a more relaxed case where the system prompt is designed and deployed to handle tasks and queries within one specific domain [^1] Unless otherwise stated, we report results with the Domain setup..
@@
Default 45.1 68.9 64.0 59.9 17.5 51.1
SPRIG 45.4 69.3 65.3 64.7 17.7 52.5
MetaSPO 45.6 71.4 67.3 67.2 19.9 54.3
Figure 6: Efficiency for test-time adaptation as a function of optimization iterations (left) and data quantity (right). The results are averaged over the Review Analysis and Reasoning domains. The gray dashed lines indicate the final performance of the Default baseline.
@@ 4.2 Experimental Results and Analyses
We first report the performance of optimized system prompts on the unseen generalization scenario, where examples for target tasks are not available for prompt optimization; thus, the user prompts are not optimized on them, and, for evaluation, (ten) user prompts on each task are obtained via LLM prompting (see A.5 for details). As shown in Table 1, we then observe that MetaSPO consistently outperforms all baselines across both the Global and Domain system prompt optimization settings. To see whether the optimized system prompt contributes to the performance gain when coupled with diverse user prompts, we visualize the performance of (randomly sampled 20%) user prompts using the system prompt optimized from MetaSPO, compared to the default system prompt in Figure 5 (see LABEL:scatter_domains for the break-down results by domain). From this, we observe that 85.0% of the user prompts exhibit improved performance with MetaSPO, indicating that it effectively and robustly enhances the performance across a broad range of user prompts.
@@
We hypothesize that if the target tasks (used for evaluation) are more similar to the source tasks (used for meta-learning), the system prompt optimized from source tasks is more effective for target tasks. To confirm this, we measure the similarity of examples between source and target tasks using either Bag-of-Words rank correlation [21] or cosine similarity over the embedding space with the language model [33], then measure the Pearson correlation of the similarity with its corresponding performance improvement (over the Default system prompt). The results in Figure 5 show a positive correlation between the source-target task similarity and the improvement, with a Pearson correlation coefficient of +0.56. Also, the cosine similarity result in LABEL:embedding_cosine_similarity shows a positive correlation with performance improvement. Yet, more interestingly, we observe that MetaSPO remains effective even for low-similarity tasks, yielding performance gains. These results demonstrate that, while including more source tasks that are close to target tasks is beneficial, MetaSPO enables learning generalizable knowledge that is useful across diverse tasks.
@@
Table 3: Results with different LLMs for MetaSPO, where Cross Model refers to prompts optimized with Llama3.2 (3B) and applied to other models. Results are averaged over Review Analysis and Reasoning domains. Numbers in bold indicate the highest, followed by underline.
@@
[0.5pt/1pt] Cross Model 70.1 68.3 78.3
Table 4: Variations of MetaSPO, where we use different compositions of optimization methods within our meta-learning framework. We report the average performance score for each domain. The highest score in each domain is highlighted in bold, while the second-highest is underlined.
@@
We conduct an auxiliary analysis to examine whether MetaSPO is versatile with other LLMs (for response generation) and whether the optimized system prompt from one LLM can be generalizable to other LLMs. For both experiments, we consider the following LLMs: Llama 3.1 (8B), Qwen 2.5 (7B), and GPT-4o mini. We then report the results in subsection 4.2, and, from this, we observe that MetaSPO demonstrates its robustness and generalizability. Specifically, the system prompt optimized by MetaSPO is superior to other baselines regardless of the underlying LLMs used as the base model. Furthermore, in the Cross Model, the system prompt optimized for Llama 3.2 (3B) demonstrates strong generalization capabilities when applied to LLMs other than it without requiring additional optimization. Overall, these results confirm that MetaSPO is effective for diverse LLMs, producing system prompts that not only perform well within the LLM they were optimized for but also maintain high performance across different LLMs. Lastly, we further extend our analysis of MetaSPO with different optimizer LLMs and show its robustness, provided in LABEL:other_optimizers.
@@
We present examples for the optimized system prompts in LABEL:qualitative, from which we observe that they typically provide a more specific role to the LLM beyond the helpful assistant (e.g., you are a knowledgeable and analytical assistant designed to process and analyze information across a broad range of topics) and include detailed guidelines or objectives on top of it.
@@ 5 Conclusion
In this paper, we introduced the novel problem of bilevel system prompt optimization, which differs from existing prompt optimization work (that primarily targets optimization of task-specific user prompts) and instead focuses on designing system prompts that are robust to diverse user prompts and transferable to unseen tasks and domains. To address this problem, we then proposed MetaSPO, a meta-learning framework that optimizes system prompts over a wide range of user prompts and tasks, while iteratively refining the user prompts and using those optimized user prompts during system prompt optimization to ensure effective generalization of the system prompt to various (optimized) user prompts. We extensively validated the proposed MetaSPO framework on 14 unseen datasets across 5 distinct domains, on which we demonstrated that it consistently outperforms baselines in both unseen generalization and test-time adaptation scenarios. We believe our work establishes a significant step forward in enhancing the robustness and adaptability of LLMs, enabling the use of optimized system prompts designed to generalize across diverse tasks and domains as well as LLMs.
Limitation and Societal Impacts
Limitation
Despite the promising results of MetaSPO, our work has some interesting avenues for future work. Specifically, while we demonstrate that MetaSPO works effectively even with a small open-source optimizer (Table LABEL:tab:Optimizer_Ablation), its performance depends on the capability of the optimizer LLM. We note that this is indeed a common challenge shared across many prompt optimization studies, and to tackle this, developing approaches to maintain strong performance with smaller (open-sourced) LLMs for optimization (such as with model distillation) would be an exciting avenue for future work.
Societal Impacts
We strongly believe that our MetaSPO (designed to formulate system prompts that are generalizable across diverse tasks and domains, as well as user prompts) has huge potential for broad real-world applications. In the meantime, as the performance of MetaSPO is still far from perfect, for the mission-critical domains (such as biomedical), it should be used carefully. In addition to this, there is a chance that MetaSPO is misused to steer the behavior of LLMs in harmful ways, and while this vulnerability is not unique to our approach but a common challenge faced by existing prompt optimization methods, an additional safeguard for it may be needed.
Acknowledgements
This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST), No.RS-2022-II220713, Meta-learning Applicable to Real-world Problems ), National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), a grant of the Korea Machine Learning Ledger Orchestration for Drug Discovery Project (K-MELLODDY), funded by the Ministry of Health & Welfare and Ministry of Science and ICT, Republic of Korea (No. RS2024-00460870) , Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. (No. RS-2024-00469482 & RS-2024-00509279) , and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT(MSIT, Korea) & Gwangju Metropolitan City, and i-Scream Media.
@@ A.2 Additional Implementation Details
We now provide the additional implementation details for other optimization methods included in our experiments. Regarding ProTeGi [27] (presented in Table 6), it performs six iterations with a beam size of three. For the Outer Loop, we perform six iterations, which is twice the number of iterations used in MetaSPO. This adjustment ensures an equivalent total number of iterations, as the Outer Loop method lacks an inner loop process. In the case of MetaSPO w/ APE and MetaSPO w/EVO (that are operationalized without the prompt analysis step), we generate 18 new candidate prompts through resampling, crossover, and mutation. For SPRIG, we conduct the experiments using the implementation provided in its official repository referenced in the original paper [43]. We iterate SPRIG three times to ensure a comparable amount of computation to the proposed MetaSPO (See LABEL:appendix:computational_cost for details). Experiments are primarily conducted using an NVIDIA A5000 GPU.
@@ A.3 Algorithm of MetaSPO
0:  Task distribution $\mathcal{T}$, Initial system prompt $\bm{s}$}, Number of iterations NN
@@ A.5 Generation of User Prompts for Unseen Generalization
In the unseen generalization setting, we evaluate whether the system prompt is effective with user prompts not optimized for target tasks. To generate these unoptimized user prompts, we provide GPT-4o mini with ten input-output example pairs, generating the coarse user prompts following Zhou et al. [45] (please see Table 11 for the detailed meta prompt used to generate ten different user prompts). After that, the answer format prompts (e.g., At the end, present your answer in <answer>yes</answer> or <answer>no</answer>.) are added to those generated user prompts. Note that the examples of the generated user prompts for each target task are provided in LABEL:tab:example_up_unseen_gen.
Table 11: Meta Prompt for generating ten user prompts for the unseen generalization scenario.